The ‘Kluwer’ LaTeX Style File: An example opening environment^{∗}

Basic Instructions

A. Thor1 (thor@wkap.nl) and B. Thor2^{†}

KAP

C. Thor3

Abstract. This document describes how to format a paper with kluwer.cls, the Kluwer classfile for journal submissions.

Keywords: LaTeX, Kluwer, Journal submissions

Abbreviations: KAP – Kluwer Academic Publishers; compuscript – Electronically submitted article

JEL codes: D24, L60, 047

Nomenclature: KAP – Kluwer Academic Publishers; compuscript – Electronically submitted article

De Kluwer LaTeX stylefile; aanwijzingen voor auteurs

To Jim

What can’t be done with TeX isn’t worth doing.

1. A heading

Some text.

∗ Donald Knuth † Partially support...

\begin{matrix}c \\ ◯ \end{matrix} 2000 Kluwer Academic Publishers. Printed in the Netherlands.

2. A heading

Some text.

3. A heading

Some text.

Address for Offprints: Kluwer Prepress Department P.O. Box 990 3300 AZ Dordrecht The Netherlands